OpenStudy (anonymous):

What does 4 c = d for c

OpenStudy (anonymous):

4c=d 4c/4 = d/4 c = d/4

OpenStudy (anonymous):

Are you solving equations with only/mostly variables? I just learned this like 2 days ago

OpenStudy (anonymous):

Yea I'm new to this and idk what to do

OpenStudy (anonymous):

And the equations are getting bigger like n-6m=8 for n

OpenStudy (anonymous):

And what does q for p and 8 for n mean? :)

OpenStudy (anonymous):

You still there? I'm sorry I was doing my homework